What are the next three terms of the geometric sequence 5, 15, 45, ? Thank you
Found 2 solutions by benazir.sj@gmail.com, ej_03:
Answer by benazir.sj@gmail.com(109)

the general form of geometric sequence is
a,ar,ar^2,ar^3.........
where a is the first number
r is the common ratio
given 5, 15, 45
a=5 and r =3
a =5
ar =5*3=15
ar^2=5*3^2
=5*9 =45
ar^3 =5*3^3
=5*27 =135
ar^4 =5*3^4
=5*81=405
ar^5=5*3^5
=5*243=1215
therefore the geometric sequence are
5,15,45,135,405,1215...



Answer by ej_03(8)

135, 405,1215.... where each given sequence of number are multiplied by 3...